# User Simulation for Evaluating Information Access Systems

### v1

@@ 1.5 Aims and Organization
With the emergence of various information access systems exhibiting increasing complexity, there is a critical need for sound and scalable means of automatic evaluation. Simulation has the potential to offer a solution here. It has attracted attention from multiple angles and much progress has been made in the last decade or so. Relevant research work, however, has been scattered in multiple research communities, including information retrieval, recommender systems, dialogue systems, and user modeling. This book aims to synthesize that research into a coherent framework. Given the substantial amount of work performed within the context of information access systems, this is where our main focus will lie.
Specifically, our main objective is to discuss how simulation may be employed to undertake evaluation of information access systems in order to (1) estimate how well they will perform under various circumstances, and (2) analyze how performance changes under different conditions and user behaviours. However, we attempt to make our discussions as generic as possible, such that those working on other types of interactive systems, or applications of assistive AI, would also find it useful.
@@
The intended audience of this book includes both researchers, who wish to further the development of simulation-based evaluation methods, as well as industry practitioners, who are interested in employing these techniques in operational settings.
@@
The rest of this book is organized as follows. We start in Chapter 2 by providing a background on the development of simulation techniques within different research communities. Following this, Chapter 3 gives an overview of how user simulation has been employed in the past. Chapter 4 introduces a conceptual framework for generally modeling interactions between a user and a system. Next, in Chapter 5 we discuss decision-making and cognitive processes of users, followed by the mathematical framework we will employ in subsequent chapters to model these. We present simulation techniques for search engines and recommender systems in Chapter 6 and for conversational assistants in Chapter 7. Finally, Chapter 8 concludes the book by highlighting open issues and possible future research directions.
@@
2.1 Information Retrieval and Recommender Systems
Both search and recommendation address the problem of providing users with items that are estimated to be relevant to the user’s information need, preferences, and/or context, often presented as a ranked list. As they share a lot of commonalities in terms of evaluation methodology, we discuss these two tasks together.
The information retrieval community has made significant impact on evaluation methodology and developed many widely used measures such as Precision, Recall, F1, Average Precision, and nDCG [Sanderson:2010:FnTIR]. Due to the empirical nature of information retrieval, researchers in the IR field realized the importance of reproducible quantitative evaluation of an IR system as early as in 1960s when the Cranfield evaluation methodology was developed and basic measures such as Precision and Recall were proposed. The test collection-based evaluation methodology has since become the standard, not only for evaluating IR tasks, but also virtually all other empirical tasks in AI, including problems in natural language processing, machine learning, and computer vision.
@@ 2.1 Information Retrieval and Recommender Systems
From a simulation perspective, the Cranfield evaluation methodology can be regarded as evaluation based on simulating overly simplified users without consideration of users’ interactions. Given the amount of interactive support today’s search engines provide for users, it is clear that the Cranfield methodology is no longer adequate in such highly interactive settings. Recognizing this, industrial research labs often use A/B tests and interleaved comparisons to compare alternative methods or algorithms [Hofmann:2016:FnTIR]. Observations made on real users’ behaviour in their natural environment, if interpreted correctly, can yield the most accurate measurement of overall system effectiveness. However, a major limitation is that online experiments cannot be reproduced. As will be discussed later in the book, user simulation techniques address these limitations, and generalize the Cranfield evaluation methodology to enable evaluation of not only interactive search systems, but also recommender systems, and assistive AI systems in general.
Simulation has a long history in IR. The first study that used simulation for the evaluation of IR systems was by Cooper:1973:ISR. [^1] Work on simulation in IR dates even further back, but the focus was different. For example, Blunt:1965:techreport presented a simulation approach to measure the effect of various system configurations on cost and on response time. Interestingly, costs were not only associated with equipment but also with personnel. They generated synthetic queries and documents in order to analyze the effect of changes in query characteristics on the number of documents retrieved. Note that the relevance of documents is not considered, but this is acknowledged as one of the main limitations of the study. SparkJones:1979:JD and Harman:1992:SIGIR studied the effectiveness of relevance feedback with the help of simulation.
It was in the first decade of the 2000s when simulation got traction, mostly within researchers working on interactive information retrieval. For instance, Leuski:2000:CIKM studied relevance feedback using neural networks and reinforcement learning, and Jordan:2006:JCDL and Azzopardi:2006:SIGIR generated synthetic queries. The SIGIR 2010 workshop on the Simulation of Interaction (SimInt) [Azzopardi:2011:SIGIRForum] represents a milestone event. Some of the most influential work on simulation around this time focused on query formulation [Baskaya:2012:SIGIR], relevance feedback [Keskustalo:2008:IR], scanning/examination/stopping behaviour [Turpin:2009:SIGIR, Baskaya:2013:CIKM, Maxwell:2015:CIKM], and the relationship between cost and effort [Azzopardi:2009:SIGIR, Azzopardi:2011:SIGIR, Baskaya:2012:SIGIR]. This was followed by a relatively quiet period, until a renewed interest in simulation has been triggered, in a large part due to recent research on conversational information access systems (cf. Section 2.1.3). This renewed interest was also marked by the Simulation for IR Evaluation (Sim4IR) workshop at SIGIR’21 [Balog:2022:SIGIRForum].
It is important to realize that elements of user simulation have been prevalent in the field of IR, even if they have usually not been called as such—virtually all evaluation measures assume some sort of user model, and thus may be viewed as naive simulators (cf. Section 4.1).
2.1.1 Interactive IR
@@
Most recently, there has been progress in formal models for IIR, including the probability ranking principle for IIR [Fuhr:2008:IRJ], economic models [Azzopardi:2011:SIGIR, Azzopardi:2014:SIGIR], the interface card model [Zhang:2015:SIGIR], and the game theoretic framework for IR [Zhai:2016:MC].
@@
Compared to the great deal of work on static test collections [Sanderson:2010:FnTIR], there has been limited research on test collections that model dynamic user interactions. Attempts to go beyond static test collections include the TREC Query track, with multiple queries for the same topic [Buckley:1999:TREC], the TREC Web track, with multi-faceted relevance judgments [Clarke:2011:TREC], and the TREC Session track, with actual user interactions in a session history [Carterette:2016:SIGIR].
2.1.2 Recommender Systems
Early work on recommender systems focused on the problem of rating prediction: predicting a given user’s rating for a given item based on past ratings or other information. It was then realized that this is an oversimplified abstraction of the actual user task. Users are presented with a list of recommended items and would first select which items to . This top-k recommendation task has since become the standard, adopting IR evaluation measures. However, a major difference from IR, from an evaluation perspective, is that users provide ratings themselves; there is no need to collect relevance assessments by third-party raters. Also, unlike in search, where users have the initiative, recommendations are driven by system initiative. What items are shown to users as suggestions influences what choices they can make. These can lead to various biases.
Recently, simulation has been used for measuring potential long-term effects of interactions between users and the recommender system [McInerney:2021:RecSys, Yao:2021:arXiv, Stavinova:2022:arXiv]. The SimuRec workshop held at RecSys’21 [Ekstrand:2021:RecSys] is further evidence that the need for simulation techniques has been recognized within the RecSys community.
2.1.3 Conversational Search and Recommendation
@@
Conversational search has been recognized as a frontier of strategic importance within IR [Anand:2020:Dagstuhl, Culpepper:2018:SIGIRForum]. Similarly, in the RecSys community, conversational interfaces to recommendations have gained research attention [Jannach:2021:CSUR, Gao:2021:AI]. Given the close relationship between search and recommendation, the delineation of the two tasks in a conversational setting is even more challenging. Therefore, the term conversational information access [Balog:2021:DESIRES] or conversational information seeking [Zamani:2023:arXiv] is used to represent a broader scope that encompasses the different conversational goals or subdomains (search, recommendation, and also question answering). The advent of conversational information access systems accelerated the need for offline evaluation methods that consider user interactions [Lipani:2021:TOIS, Salle:2021:ECIR, Zhang:2020:KDD, Salle:2021:ECIR, Sekulic:2022:WSDM, Owoicho:2023:SIGIR], which has triggered a renewed interest in user simulation techniques [Balog:2022:SIGIRForum].
@@ 2.2 Dialogue Systems
One of the most important and fundamental ideas from dialogue systems research is modeling human-computer dialogue formally as a Markov Decision Process (MDP) [Levin:2000:TSAP, Young:1999:PTRS]. Crucially, mapping dialogue onto the MDP model requires a well-defined representation of the state space and action set. For slot-filling dialogues, a common approach is to base the states on the various slots that need to be filled and the actions on the intents of the participants. Dialogue policy learning has been approached using two different forms of reinforcement learning: model-based and simulation-based [Schatzmann:2006:KER]. The model-based approach ignores actual user responses and considers only system actions to estimate transition probabilities directly from training data [Singh:2022:JAIR, Walker:1998:COLING]. However, the vast space of possible dialogue states and strategies cannot be sufficiently explored from a fixed corpus to obtain reliable estimates; also, new strategies cannot be explored as they are not guaranteed to be present in the training data [Schatzmann:2006:KER]. Instead, a dynamic environment is needed that allows for a large number of “trial-and-error” interactions. According to the simulation-based approach, a user model is trained based on the available dialogue corpus and typically using some form of supervised learning. Then, reinforcement learning techniques can be used for finding an optimal dialogue policy based on interactions with a simulated user. This way, any number of training episodes can be generated in order to explore the space of possible strategies, including those that are not present in the original training data. Consequently, simulation-based learning has become the predominant form of learning dialogue strategies [Schatzmann:2006:KER, Young:2010:CSL].

### v2

@@ 1.5 Aims and Organization
With the emergence of various information access systems exhibiting increasing complexity, there is a critical need for sound and scalable means of automatic evaluation. Simulation has the potential to offer a solution here. It has attracted attention from multiple angles and much progress has been made in the last decade. Relevant research work, however, has been scattered in multiple research communities, including information retrieval, recommender systems, dialogue systems, and user modeling. This book aims to synthesize that research into a coherent framework. Given the substantial amount of work performed within the context of information access systems, this is where our main focus will lie.
Specifically, our main objective is to discuss how simulation may be employed to undertake evaluation of information access systems in order to (1) estimate how well they will perform under various circumstances, and (2) analyze how performance changes under different conditions and user behaviours. However, we attempt to make our discussions as generic as possible, such that those working on other types of interactive systems, or applications of assistive AI, would also find it useful. Specifically, whenever possible, we would attempt to lay out general conceptual frameworks, discuss general challenges, and extract general ideas from specific research work, which we hope to be broadly useful to more readers as well as provide a stable meaningful high-level structure where future research work can be naturally incorporated and discussed. Our emphasis on the generality of discussion, however, means that our treatment of any specific research work is inevitably brief. Detailed information can be conveniently found in the numerous research papers cited throughout this book. When selecting specific work to elaborate, we have also chosen to focus more on representative work that is useful for illustrating major ideas instead of having an even coverage of all the work. Due to the interdisciplinary nature of the topic and quick growth of research, the references cited in our book are inevitably incomplete. Considering this limitation and anticipating the rapid progress of research in this area in the future, we have created a website (https://usersim.ai/) for an envisioned broad interdisciplinary research community on user simulation. This platform aims to foster collaboration among researchers from diverse fields, enabling them to collectively maintain a repository of up-to-date and relevant references over time.
@@
The main intended audience of this book includes both researchers, who wish to further the research and development of simulation-based evaluation methods, as well as industry practitioners, who are interested in employing these techniques in operational settings. Considering the fact that user simulation is broadly connected with multiple fields (see a more detailed discussion of this in Chapter 9), we also hope this book will be broadly useful to an audience beyond those interested in using user simulation for evaluation.
@@
The rest of this book is organized as follows. We start in Chapter 2 by providing a background on the development of simulation techniques within different research communities. Following this, Chapter 3 gives an overview of how user simulation has been employed in the past. Chapter 4 introduces a conceptual framework for generally modeling interactions between a user and a system and evaluating any interactive system using user simulation. Next, in Chapter 5, we discuss decision-making and cognitive processes of users, followed by the mathematical framework we will employ in subsequent chapters to model these. We present simulation techniques for search engines and recommender systems in Chapter 6 and for conversational assistants in Chapter 7. Chapter 8 focuses on applying user simulation in practice, covering issues related to configuring, validating, and building simulators. Chapter 9 broadens the perspective on user simulation as an interdisciplinary research area intersecting with various fields beyond computer science, ultimately contributing to the progress towards AGI. Finally, Chapter 10 concludes the book by highlighting open issues and possible future research directions.
@@
2.1 Information Retrieval
@@
Due to the empirical nature of information retrieval, researchers in the IR field realized the importance of reproducible quantitative evaluation of an IR system through reusable test collections as early as in 1960s when the Cranfield evaluation methodology was developed [Cleverdon:1968:Cranfield]. However, given the amount of interactive support today’s search engines provide, the static nature of test collections makes it challenging to apply this methodology in such highly interactive settings. Recognizing this, industrial research labs often use A/B tests [Hofmann:2016:FnTIR], where the idea is to interleave the results from alternative methods or algorithms, and observations made on real users’ behaviour in their natural environment, if interpreted correctly, could then be analyzed to assess the overall effectiveness of the systems or methods to be compared. However, a major limitation is that online experiments cannot be reproduced. User simulation techniques address these limitations, and generalize the Cranfield evaluation methodology to enable evaluation of not only interactive search systems, but also recommender systems, and assistive AI systems in general.
Below, we briefly introduce test collection-based evaluation (Section 2.1.1), followed by research on interactive IR systems (Section 2.1.2) and conversational information access systems (Section 2.1.3). A high-level historical account on the use of simulation in IR is presented in Section 2.1.4.
2.1.1 Test Collection-Based Evaluation
Test collection-based evaluation [Sanderson:2010:FnTIR] is the primary evaluation methodology used for assessing the utility of information access systems. The methodology was originally developed by information retrieval researchers in the 1960s to quantitatively evaluate information retrieval systems [Cleverdon:1968:Cranfield]. The methodology is often referred to as the Cranfield evaluation methodology. [^1] https://en.wikipedia.org/wiki/Cranfield_experiments A test collection includes three components: (1) a set of documents, (2) a set of queries, and (3) relevance judgments of documents in the collection with respect to each query made by human assessors. A retrieval system would be evaluated by running the system to process each query on the document set to produce retrieval results. The relevance judgments could then be used to compute various measures for assessment of the quality of the retrieval results. For example, Precision and Recall are two basic measures proposed for evaluating a set of retrieved documents, with Precision indicating the percent of retrieved documents that are relevant according to relevance judgments, whereas Recall the percent of all the relevant documents in the collection that have been retrieved [Cleverdon:1968:Cranfield]. Other measures such as F1 [VanRijsbergen:1979:book] can be used to give a single number summary of precision and recall (harmonic mean of Precision and Recall). A ranked list can be evaluated by using Average Precision [VanRijsbergen:1979:book, Harman:1995:IPM] and Normalized Cumulative Discounted Gain (NDCG) [Jarvelin:2002:TOIS]. The test collection-based evaluation methodology has since become the standard and been widely used in not only in the IR community but also many other communities for evaluation of virtually all kinds of empirical tasks, notably those in natural language processing, computer vision, and machine learning, and measures such as Precision, Recall, F1, Average Precision, and NDCG remain the main measures used today for evaluating IR systems, as well as many other intelligent systems in general [Burnell:2023:Science].
The test collection-based evaluation can be regarded a simple form of evaluation using simulation where the simulated users are implicitly specified with many unrealistic assumptions made about the users. Despite that those strong assumptions made are not valid, the methodology has been shown to be reasonable and even reliable for relative comparison of retrieval systems based on studies in the context of TREC [Voorhees:2005:book]. Most of the discussions in our book are about how to generalize the test collection-based evaluation method by explicitly specifying the users to be simulated, simulating various kinds of user actions, and making more realistic user simulation in general. Another dimension of extension is to introduce general frameworks for defining user simulation and evaluation measures that can be applied across a range of interactive information access systems, including search engines, recommender systems, and conversational assistants.
2.1.2 Interactive Information Retrieval
@@
Over the years, several notable advancements have been made in formal models for IIR, including the probability ranking principle for IIR [Fuhr:2008:IRJ], economic models [Azzopardi:2011:SIGIR, Azzopardi:2014:SIGIR], the interface card model [Zhang:2015:SIGIR], and the game theoretic framework for IR [Zhai:2016:MC].
@@
Compared to the great deal of work on static test collections [Sanderson:2010:FnTIR], there has been limited research on test collections that model dynamic user interactions, which might include, e.g., a user’s query clickthrough, query reformulation, and multiple queries in a whole session. However, several attempts have been made to go beyond static test collections, including the TREC Interactive track, where interactive query formulation was studied [Over:2001:IPM], the TREC Query track, with multiple queries for the same topic [Buckley:1999:TREC], the TREC Web track, with multi-faceted relevance judgments [Clarke:2011:TREC], and the TREC Session track, with actual user interactions in a session history [Carterette:2016:SIGIR]. In addition to TREC, evaluation of interactive retrieval tasks has also been studied in the context of CLEF [Ferro:2019:Book] and NTCIR [Sakai:2021:Book].
2.1.3 Conversational Information Access
Even though interactions between a user and a traditional search/recommender system are a form of conversation, the users of those systems have access to only a limited set of actions allowed by a system’s user interface (querying, clicking, rating, etc.). In contrast, in a truly conversational setting, the user can freely express their request and feedback in natural language. Since user intents are no longer captured as explicit actions, such as querying or clicks, they need to be inferred from free text, which requires more advanced natural language understanding capabilities. Even in seemingly straightforward cases, such as yes/no questions, humans might answer indirectly. For example, when asked about their interest in an evening activity, the user might respond with “I’d prefer to go to bed” instead of simply saying “no” [Louis:2020:EMNLP]. On the other hand, there is a much richer possibility for describing intent and providing feedback to the system via natural language. Overall, conversational systems can reduce effort for the user from a cognitive perspective, but there is also more potential for frustration [Luger:2016:CHI], as the user has less direct control over the behaviour of the system.
@@ 2.1.3 Conversational Information Access
Conversational search has been recognized as a frontier of strategic importance within IR [Anand:2020:Dagstuhl, Culpepper:2018:SIGIRForum]. Similarly, in the RecSys community, conversational interfaces to recommendations have gained research attention [Jannach:2021:CSUR, Gao:2021:AI]. Given the close relationship between search and recommendation, the delineation of the two tasks in a conversational setting is even more challenging. Therefore, the term conversational information access [Balog:2021:DESIRES] or conversational information seeking [Zamani:2023:FnTIR] is used to represent a broader scope that encompasses the different conversational goals or subdomains (search, recommendation, and also question answering).
Another crucial factor that sets conversational information access apart from traditional search and recommender systems is the notion of initiative. The interaction in traditional search systems is driven by the user, i.e., the user issues a request (query) and the system responds with a ranked list of results, also known as the query-response paradigm. In recommender systems, conversely, the system takes initiative by offering suggestions to the user based on their history. Mixed-initiative refers to a conversational paradigm where the user and system both actively participate in addressing the user’s information need. Mixed-initiative has been identified as a critical component of conversational information access systems [Radlinski:2017:CHIIR, Trippas:2018:CHIIR, Wadhwa:2021:DESIRES]. It enables systems to engage with users in natural language, for example, to ask clarification questions, solicit feedback, or to learn about the user’s preferences in specific contexts. Despite the great potential of such system-initiated questions, evaluation challenges are a bottleneck to progress: since the next system response is conditioned on the user’s answer, and the user is not constrained to a predefined set of possible answers, it is not possible to create reusable offline test collections for end-to-end evaluation [Balog:2021:DESIRES].
The advent of conversational information access systems accelerated the need for offline evaluation methods that consider user interactions and has triggered interest in user simulation techniques [Balog:2021:DESIRES, Lipani:2021:TOIS, Salle:2021:ECIR, Zhang:2020:KDD, Salle:2021:ECIR, Sekulic:2022:WSDM, Owoicho:2023:SIGIR].
2.1.4 Simulation in Information Retrieval
The general idea of simulation was explored in IR many decades ago. For example, Blunt:1965:techreport presented a simulation approach to measure the effect of various system configurations on cost and on response time. Interestingly, costs were not only associated with equipment but also with personnel. Cooper:1973:ISR used simulation to generate synthetic queries and documents in order to analyze the effect of changes in query characteristics on the number of documents retrieved, but the relevance of documents was not considered, a main limitations acknowledged in the study. Griffiths:1976:PhDThesis introduced a general simulation framework for IR systems and proposed to estimate its parameters from empirical data. Tague:1980:SIGIR employed user simulation in bibliographic retrieval systems. SparkJones:1979:JD and Harman:1992:SIGIR studied the effectiveness of relevance feedback with the help of simulation. Gordon:1990:JASIS measured the effectiveness of document retrieval systems using of simulated queries.
It was in the first decade of the 2000s when simulation gained significant traction, mostly within researchers working on interactive information retrieval. Simulation has been used for studying relevance feedback [Leuski:2000:CIKM, White:2004:ECIR, White:2005:TOIS, Keskustalo:2008:IR], information filtering [Mostafa:2003:IR], and online vertical selection [Diaz:2009:SIGIR]. Another line of work has focused synthetic query generation [Jordan:2006:JCDL, Azzopardi:2006:SIGIR], The SIGIR 2010 workshop on the Simulation of Interaction (SimInt) [Azzopardi:2011:SIGIRForum] represents a milestone event. Some of the most influential work on simulation around this time focused on query formulation [Baskaya:2012:SIGIR], scanning/examination/stopping behaviour [Turpin:2009:SIGIR, Baskaya:2013:CIKM, Maxwell:2015:CIKM], and the relationship between cost and effort [Azzopardi:2009:SIGIR, Azzopardi:2011:SIGIR, Baskaya:2012:SIGIR]. This was followed by a relatively quiet period, until a renewed interest in simulation has been triggered, in a large part due to recent research on conversational information access systems (cf. Section 2.1.3). This renewed interest was also marked by the Simulation for IR Evaluation (Sim4IR) workshop at SIGIR’21 [Balog:2022:SIGIRForum].
It is important to realize that elements of user simulation have been prevalent in the field of IR, even if they have usually not been called as such—virtually all evaluation measures assume some sort of user model, and thus may be viewed as naive simulators (cf. Section 4.1). While not treated in this book, a broader view of simulation could go beyond simulating the users themselves to simulate their whole task environment, particularly the simulation of test collections [Hawking:2020:Book].
2.2 Recommender Systems
Both search and recommendation address the problem of providing users with items that are estimated to be relevant to the user’s information need, preferences, and/or context, often presented as a ranked list.
Early work on recommender systems focused on the problem of rating prediction: predicting a given user’s rating for a given item based [Steck:2013:RecSys]. It was then realized that this is an oversimplified abstraction of the actual user task. Users are presented with a list of recommended items and must then decide which items to explore further, taking into account various factors such as personal preferences, contextual information, and item characteristics. This top-$k$ recommendation task has since become the standard, adopting IR evaluation measures [Cremonesi:2010:RecSys, Bellogin:2011:RecSys]. However, a major difference from IR, from an evaluation perspective, is that users provide ratings themselves; there is no need to collect relevance assessments by third-party raters. Also, unlike in search, where users have the initiative, recommendations are driven by system initiative. What items are shown to users as suggestions influences what choices they can make. These can lead to various biases.
Over the years, different types of interactive recommender systems have been proposed, including critiquing-based systems [Chen:2012:UMUAI], user-controlled recommenders [Jannach:2017:ECWeb], visual recommenders [Kunkel:2017:IUI, Verbert:2016:TOIS], and conversational recommenders [Gao:2021:AI, Jannach:2021:CSUR]. The evaluation of such interactive systems necessitates some form of human involvement, i.e., either real or simulated users.
One of the earliest works of using simulation for evaluation is by Dzyabura:2013:RecSys, who propose mechanisms to combine search and recommendations and demonstrate the benefits of such approach via simulations. Recently, simulation has been used for measuring potential long-term effects of interactions between users and the recommender system [McInerney:2021:RecSys, Yao:2021:arXiv, Stavinova:2022:arXiv, Hazrati:2022:IPM] as well as for developing evaluation measures specific to recommender systems that employ carousel-based result presentation [Felicioni:2021:UMAP, Dacrema:2022:FBD, Rahdari:2024:TORS]. The SimuRec workshop held at RecSys’21 [Ekstrand:2021:RecSys] is further evidence that the need for simulation techniques has been recognized within the RecSys community. Overall, “although the volume of simulation-based research in the recommender system context is gradually increasing, simulations are most frequently used to explore the impact of recommender systems on various aspects of user behaviour rather than to assess their performance and effectiveness in a comparative way” [Rahdari:2024:TORS].
@@ 2.3 Dialogue Systems
One of the most important and fundamental ideas from dialogue systems research is modeling human-computer dialogue formally as a Markov Decision Process (MDP) [Levin:2000:TSAP, Young:1999:PTRS], which we will further discuss as a general formal framework for user simulation in the book (see Section 5.3). Crucially, mapping dialogue onto the MDP model requires a well-defined representation of the state space and action set. For slot-filling dialogues, a common approach is to base the states on the various slots that need to be filled and the actions on the intents of the participants. Dialogue policy learning has been approached using two different forms of reinforcement learning: model-based and simulation-based [Schatzmann:2006:KER]. The model-based approach ignores actual user responses and considers only system actions to estimate transition probabilities directly from training data [Singh:2022:JAIR, Walker:1998:COLING]. However, the vast space of possible dialogue states and strategies cannot be sufficiently explored from a fixed corpus to obtain reliable estimates; also, new strategies cannot be explored as they are not guaranteed to be present in the training data [Schatzmann:2006:KER]. Instead, a dynamic environment is needed that allows for a large number of “trial-and-error” interactions. According to the simulation-based approach, a user model is trained based on the available dialogue corpus and typically using some form of supervised learning. Then, reinforcement learning techniques can be used for finding an optimal dialogue policy based on interactions with a simulated user. This way, any number of training episodes can be generated in order to explore the space of possible strategies, including those that are not present in the original training data. Consequently, simulation-based learning has become the predominant form of learning dialogue strategies [Schatzmann:2006:KER, Young:2010:CSL].